OpenStudy (anonymous):

Find the area of the triangle formed by the x-axis and the lines tangent and normal to the curve f(x)=7x-3x^2 at the point (2,2)

OpenStudy (anonymous):

slope of tangent = f'(x) = 7 - 6x = 7 -12 = -5 at (2,2) slope of the normal = 1/5 equation of tangent is y - 2 =-5(x-2) and of normal is y - 2 = 1/5(x-2) intercept on x-axis of tangent is 2.4 intercept on x-axis of normal is 5 so the length of base of the triangle is 5-2.4 = 2.6 area of triangle = 1.3 * 2 = 2.6

OpenStudy (anonymous):

Thanks Jimmyrep! just saw this now. will go through it to make certain I understand.